40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 275 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  3044   Thu Jun 3 14:01:51 2010 kiwamuUpdateIOOmode measurement of new input optics

I checked the measured data of the mode profile which was taken on the last Tuesday.

For the vertical profile,

the plot shows a good agreement between the expected radius which is computed from the past measurement, and that measured on the last Tuesday.

However for the horizontal profile,

it looks like being overestimated. This disagreement may come from the interference imposed on the Gaussian spot as we've been worried. 

So I guess we should solve this issue before restarting this mode matching work.

 - The next step we should do are;

checking the effect of MMT1 on the shape of the beam spot by using a spare of MMT1

 

NOTE:

The expected curve in the attached figure were computed by using the fitted parameter listed on the entry 2984 .

In the calculation the MMT1 is placed at 1911mm away from MC3 as we measured.

And the focal length of MMT1 is set to be f=-2500mm.

Attachment 1: check_measurement_edit.png
check_measurement_edit.png
  3043   Thu Jun 3 13:14:27 2010 kiwamuUpdateIOOmode measurement of new input optics

I corrected the sketch of the new IOO. 

The sketch in the last entry was also replaced by the new one. 

http://nodus.ligo.caltech.edu:8080/40m/3029 

Quote:

Just note that MMT1 has RoC of -5m (negative!). This means that it is a lens with f=-2.5 m,

Attachment 1: inside_vac_2.png
inside_vac_2.png
  3042   Thu Jun 3 00:47:17 2010 KevinUpdatePSL2W Beam Profile of Second Reflected Beam

[Koji, Kevin]

The profile of the Innolight 2W was previously measured by measuring the reflected beam from the front surface of a W2 window (see entry). To investigate thermal effects, Rana suggested also measuring the profile of the beam reflected from the back surface of the W2.

I used the same setup and methods as were used in the first measurement. The mirror was moved so that only the beam reflected from the back surface of the W2 was reflected from the mirror. This beam was reflected from both the front of the mirror and the back of the mirror. An extra beam dump was positioned to block the reflection from the back of the mirror.

This measurement was made with 2.004 A injection current and 25.04°C laser crystal temperature.

The data was fit to w = sqrt(w0^2+lambda^2*(x-x0)^2/(pi*w0)^2) with the following results

For the horizontal beam profile:

reduced chi^2 = 5.1

x0 = (-186 ± 6) mm

w0 = (125.8 ± 1.4) µm


For the vertical beam profile:

reduced chi^2 = 14.4

x0 = (-202 ± 11) mm

w0 = (132.5 ± 2.7) µm


In the following plots, the blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

Attachment 1: profile_2nd.png
profile_2nd.png
  3041   Wed Jun 2 22:58:04 2010 KevinUpdatePSL2W Second Reflected Beam Profile

[Koji, Kevin]

The profile of the Innolight 2W was previously measured by measuring the reflected beam from the front surface of a W2 window (see entry). To investigate thermal effects, Rana suggested also measuring the profile of the beam reflected from the back surface of the W2.

I used the same setup and methods as were used in the first measurement. The mirror was moved so that only the beam reflected from the back surface of the W2 was reflected from the mirror. This beam was reflected from both the front of the mirror and the back of the mirror. An extra beam dump was positioned to block the reflection from the back of the mirror.

This measurement was made with 2.004 A injection current and 25.04°C laser crystal temperature.

The data was fit to w = sqrt(w0^2+lambda^2*(x-x0)^2/(pi*w0)^2) with the following results

For the horizontal beam profile:

reduced chi^2 = 5.1

x0 = (-186 ± 6) mm

w0 = (125.8 ± 1.4) µm

For the vertical beam profile:

reduced chi^2 = 14.4

x0 = (-202 ± 11) mm

w0 = (132.5 ± 2.7) µm

In the following plots, the blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

The differences between the beam radii for the beam reflected from the front surface and the beam reflected from the back surface are

Δw0_horizontal = (12.8 ± 1.6) µm

Δw0_vertical = (8.5 ± 2.9) µm

So the two measurements are not consistent. This suggests that the passage through the W2 altered the profile of the beam.

Attachment 1: profile_2nd.png
profile_2nd.png
  3040   Wed Jun 2 22:25:39 2010 KevinUpdatePSLLow Power 2W Beam Profile

Koji is worried about thermal lensing introducing errors to the measurement of the 2W beam profile so I measured the profile at a lower power.

I used the same setup and methods used to measure the profile at 2W (see entry). This measurement was taken with an injection current of 1.202 A and a laser crystal temperature of 25.05° C. This corresponds to approximately 600 mW (see power measurement).

The data was fit to  w = sqrt(w0^2+lambda^2*(x-x0)^2/(pi*w0)^2) with the following results

For the horizontal beam profile:

reduced chi^2 = 2.7

x0 = (-203 ± 3) mm

w0 = (151.3 ± 1.0) µm

For the vertical beam profile:

reduced chi^2 = 6.8

x0 = (-223 ± 6) mm

w0 = (167.5 ± 2.2) µm

In the following plots, the blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

The differences between the beam radii for the low power and high power measurements are

Δw0_horizontal = (38.3 ± 1.2) µm

Δw0_vertical = (43.5 ± 2.4) µm

Thus, the two measurements are not consistent. To determine if the thermal lensing is in the laser itself or due to reflection from the W2 and mirror, we should measure the beam profile again at 2W with a razor blade just before the W2 and a photodiode to measure the intensity of the reflection off of the front surface. If this measurement is consistent with the measurement made with the beam scan, this would suggest that the thermal lensing is in the laser itself and that there are no effects due to reflection from the W2 and mirror. If the measurement is not consistent, we should do the same measurement at low power to compare with the measurement described in this entry.


Attachment 1: profile_low.png
profile_low.png
  3039   Wed Jun 2 21:21:43 2010 steveUpdatePEMseismometers off of linoleum floor

Quote:

Steve for Nancy,

Seismometer interface box ac power was turned off, Guralps disconnected and moved. Ranger locked, moved and released. Nancy will describe the rest soon.

The flattened lead balls were checked for their heights by the calliper, and were all in the range of 9.50 to 9.70 mm.

The rechecking was done by using these balls between two aluminium plates and checking their levelling. When confirmed this, we proceeded to install the balls(no more balls :P ) in their place.

The Guralps were switched off by switching off the power supply to them. The ranger mass was clamped in order to be able to move it. This can be undone by rotating the transport screw counter-clockwise.

We installed the flattened lead ballsin the space made for them. The granite was then placed on it with the help of many other people in the lab.  It was lowered by hanging it on two straps held by people , and then placed in the space marked for it.

Did we then turn on the seismometers? Did we release the locking screw on the Ranger? What happened to Bat-Boy??? Since they make a good mystery I will choose to leave them out of my elog entry.

  3038   Wed Jun 2 18:36:20 2010 valeraDAQCDSNoise generators in LSP

Alex wrote a new code to implement LSP noise generator. The code is based on 64 bit random number generator from Numerical Recipes 3rd ed ch 7.1 (p 343).

Joe made two instances in the LSP model.

The attached plot shows the spectra and coherence of two generators. The incoherence is ~1/Navg - statistically consistent with no coherence.

Attachment 1: noisegenerators.pdf
noisegenerators.pdf
  3037   Wed Jun 2 18:09:32 2010 steveUpdatePEMseismometers off of linoleum floor

Steve for Nancy,

Seismometer interface box ac power was turned off, Guralps disconnected and moved. Ranger locked, moved and released. Nancy will describe the rest soon.

 

  3036   Wed Jun 2 17:34:33 2010 josephb, alex, valeraUpdateCDSCDS updates

From what I understand, Alex rewrote portions of the framebuilder and testpoint codes and then recompiled them in order to get more than 1 testpoint per front end working.   I've tested up to 5 testpoints at once so far, and it worked.

We also have a new noise component added to the RCG code.  This piece of code uses the random number generator from chapter 7.1 of Numerical Recipies Third Edition to generate uniform numbers from 0 to 1.  By placing a filter bank after it should give us sufficient flexibility in generating the necessary noise types.  We did a coherence test between two instances of this noise piece, and they looked pretty incoherent.  Valera will add a picture of it when it finishe 1000 averages to this elog.

I'm in the process of propagating the old suspension control filters to the new RCG filter banks to give us a starting point.  Tomorrow Valera and I are planning to choose a subset of the plant filters  and put them in, and then work out some initial control filters to correspond to the plant.  I also need to think about adding the anti-aliasing filters and whitening/dewhitening filters.

 

  3035   Wed Jun 2 11:28:31 2010 KojiUpdateMOPAlaser headtemp is up

Last night we stopped the air conditioning. It made HDTEMP increase.
Later we restored them and the temperature slowly recovered. I don't know why the recovery was so slow.

Quote:

Is the cooling line clogged? The chiller temp is 21C See 1 and 20 days plots

 

  3034   Wed Jun 2 11:25:16 2010 josephb,alexUpdateCDSCDS saga (aka the bad code saga)

Alex updated the awg.par file to handle all the testpoints.  Basically its very similar to the testpoint.par, but the prognum lines have to be 1 higher than the corresponding prognum in testpoint.par.  A entry looks like:

[C1-awg0]
hostname=192.168.1.2
prognum=0x31001002

After running "diag -i" and seeing some RPC number conflicts, we went into /cvs/cds/caltech/cds/target/gds/param/diag_C.conf and changed the line from

&chn * *  192.168.1.2 822087685 1

to

&chn * *  192.168.1.2 822087700 1

The number represents an RPC number.  This was conflicting with the RPC number associated with the awgtpman processes.  We then had to update the /etc/rpc file as well.  At the end we changed chnconf 822087685 to chnconf 822087700.  We then run /usr/sbin/xinetd reload

Lastly we edited the /etc/xinetd.d/chnconf file line

server_args             = /cvs/cds/caltech/target/gds/param/tpchn_C4.par /cvs/cds/caltech/target/gds/param/tpchn_C5.par

to

server_args             = /cvs/cds/caltech/target/gds/param/tpchn_C1.par /cvs/cds/caltech/target/gds/param/tpchn_C2.par /cvs/cds/caltech/target/gds/param/tpchn_C3.par /cvs/cds/caltech/target/gds/param/tpchn_C4.par /cvs/cds/caltech/target/gds/param/tpchn_C5.par /cvs/cds/caltech/target/gds/param/tpchn_C6.par /cvs/cds/caltech/target/gds/param/tpchn_C7.par /cvs/cds/caltech/target/gds/param/tpchn_C8.par /cvs/cds/caltech/target/gds/param/tpchn_C9.par

 

Alex also recompiled the frame builder code to be able to handle more than 7 front ends.  This involved tracking down a newer version of libtestpoint.so on c1iscex and moving it over to megatron, then going in and by hand adding the ability to have up to 10 front ends connected.

Alex has said he doesn't like this code and would like it to dynamically allocate properly for any number of servers rather than having a dumb hard coded limit.

Other changes he needs to make:

1) Get rid of set dcu_rate ## = 16384 type lines in the daqrc file.  That information is available from the /caltech/chans/C1LSC.ini type files which are automatically generated when you compile a model.  This means not having to go in by hand to update these in daqrc.

2) Get some awg.par and testpoint.par rules, so that these are automatically updates when you build a model.  Make it so it automatically assigns a prognum when read in rather than having to hard code them in by hand.

3)Slave the awgtpmans to a single clock running from the IO processor x00. This ensures they are all in sync.

 

 

 

  3033   Wed Jun 2 07:54:55 2010 steveUpdateMOPAlaser headtemp is up

Is the cooling line clogged? The chiller temp is 21C See 1 and 20 days plots

Attachment 1: htemp.jpg
htemp.jpg
Attachment 2: htemp20d.jpg
htemp20d.jpg
  3032   Wed Jun 2 04:27:02 2010 KojiUpdatePSL2W Beam Profile

This is what I already told to Kevin and Rana:

A direct output beam is one of the most difficult measurements for the mode profiling.
I worried about the thermal lensing.

Since most of the laser power goes through the substrate (BK7) of the W2 window, it may induce thermal deformation on the mirror surface.
An UV fused silica window may save the effect as the thermal expansion coefficient is 0.55e-6/K while BK7 has 7.5e-6.

In addition to the thermal deformation issue, the pick-off setup disables us to measure the beam widths near the laser aperture.

I rather prefer to persist on the razor blade then use the pick off between the blade and the PD.

I also confess that the description above came only from my knowledge, and not from any scientific confirmation including any calculation.
If we can confirm the evidence (or no evidence) of the lensing, it is a great addition to my experience.

Quote:

[Rana, Kiwamu, Kevin]

The Innolight 2W beam profile was measured with the beam scan. A W2-IF-1025-C-1064-45P window was used to reflect a small amount of the main beam. A 5101 VIS mirror was used to direct just the beam reflected from the front surface of the W2 down the table (the beam reflected from the back surface of the W2 hit the optic mount for the mirror). A razor blade beam dump was used to stop the main transmitted beam from the W2. The distance from the laser was measured from the front black face of the laser to the front face of the beam scan (this distance is not the beam path length but was the easiest and most accurate distance to measure). The vertical and horizontal beam widths were measured at 13.5% of the maximum intensity (each measurement was averaged over 100 samples). These widths were divided by 2 to get the vertical and horizontal radii.

The mirror was tilted so that the beam was close to parallel to the table. (The center of the beam fell by approximately 2.1 mm over the 474 mm that the measurement was made in).

The measurement was taken with an injection current of 2.004 A and a laser crystal temperature of 25.04°C.

This data was fit to w = sqrt(w0^2+lambda^2*(x-x0)^2/(pi*w0)^2) with lambda = 1064nm with the following results

For the horizontal beam profile:

reduced chi^2 = 4.0

x0 = (-138 ± 3) mm

w0 = (113.0 ± 0.7) µm

For the vertical beam profile:

reduced chi^2 = 14.9

x0 = (-125 ± 4) mm

w0 = (124.0 ± 1.0) µm

In the following plots, the blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

 

  3031   Wed Jun 2 03:50:03 2010 KojiUpdateIOOmode measurement of new input optics

Just note that MMT1 has RoC of -5m (negative!). This means that it is a lens with f=-2.5 m,

 

  3030   Wed Jun 2 03:24:22 2010 KevinUpdatePSL2W Beam Profile

[Rana, Kiwamu, Kevin]

The Innolight 2W beam profile was measured with the beam scan. A W2-IF-1025-C-1064-45P window was used to reflect a small amount of the main beam. A 5101 VIS mirror was used to direct just the beam reflected from the front surface of the W2 down the table (the beam reflected from the back surface of the W2 hit the optic mount for the mirror). A razor blade beam dump was used to stop the main transmitted beam from the W2. The distance from the laser was measured from the front black face of the laser to the front face of the beam scan (this distance is not the beam path length but was the easiest and most accurate distance to measure). The vertical and horizontal beam widths were measured at 13.5% of the maximum intensity (each measurement was averaged over 100 samples). These widths were divided by 2 to get the vertical and horizontal radii.

The mirror was tilted so that the beam was close to parallel to the table. (The center of the beam fell by approximately 2.1 mm over the 474 mm that the measurement was made in).

The measurement was taken with an injection current of 2.004 A and a laser crystal temperature of 25.04°C.

This data was fit to w = sqrt(w0^2+lambda^2*(x-x0)^2/(pi*w0)^2) with lambda = 1064nm with the following results

For the horizontal beam profile:

reduced chi^2 = 4.0

x0 = (-138 ± 3) mm

w0 = (113.0 ± 0.7) µm

For the vertical beam profile:

reduced chi^2 = 14.9

x0 = (-125 ± 4) mm

w0 = (124.0 ± 1.0) µm

In the following plots, the blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

Attachment 1: profile.png
profile.png
Attachment 2: errors.png
errors.png
Attachment 3: Layout.jpg
Layout.jpg
  3029   Wed Jun 2 01:47:28 2010 Alberto, KiwamuUpdateIOOmode measurement of new input optics

The mode profile of the new input optics was measured.

Although the distance between each optic was not exactly the same as the design because of narrow space,

we measured the profile after the curved mirror (MMT1) that Jenne and Kevin put in the last week.

 


(interference from MMT1)

Below is a sketch of the current optical path inside of the chamber.

inside_vac_2.png

 

In the beginning of this measurement, the angle between the incident and the reflection on MMT1 (denoted as theta on the sketch) was relatively big (~40deg) although MMT1 was actually made for 0deg incident.

At that time we found a spatially large interference imposed on the Gaussian beam at the beam scan. This is not good for mode measurement

This bad interference can be caused by an extra reflection from the back surface of MMT1 because the interference completely vanished by removing MMT1  .

In order to reduce the interference we decreased the angle theta as small as possible. Actually we made it less than 10deg which was our best due to narrow space. 

Now the interference got less and the spot looks better.

The picture below shows an example of the beam shape taken by using the beam scan.

Top panel represents the horizontal mode and bottom panel represents the vertical mode.

You can see some bumps caused by the interference on the horizontal mode, these bumps may lead to overestimation of the horizontal spot size .

 

beam_profile.png

 

(result)

 afterMMT1.png

 The above plot shows the result of the mode measurement.

 Here are the parameter obtained by fitting. The data is also attached as attachment:4

waist size for vertical  w0v [mm]  0.509 +/-0.0237
waist size for horizontal

w0h  [mm]

 0.537  +/- 0.0150
waist position from MMT1 for vertical  xv[m]  -2.91 +/- 0.214
waist position from MMT1 for horizontal xh[m]   -2.90 +/-  0.127


Attachment 3: MMT1_.dat.zip
  3028   Tue Jun 1 20:40:03 2010 KojiUpdatePSLnew HIGH-LOW value for PMC_TRANS

The alarm had kept crying. I reduced the LOW to be 0.90 and the LOLO to be 0.85 both in psl.db and with ezcawrite.

Quote:

We changed the HIGH/LOW values of the PMC_TRANS.

The edited file was updated on the svn.


Since the PMC_TRANSPD was replaced behind the pzt mirror (see the entry), its nominal value were reduced to something like ~1V from the previous value of ~2V.

In the medm screen C1PSL_PMC.adl the PMC_TRAN always indicated red because the value were low compared with the previous one.

We went to /cvs/cds/caltech/target/c1psl, then edited psl.db

- Here are the new parameters we set up in the file.

grecord(ai,"C1:PSL-PMC_PMCTRANSPD")
{

  field(LOW,"0.98")
  field(LOLO,"0.93")
  field(HIGH,"1.15")
  field(HIHI,"1.3")

}

- - - -

These values are based on ~4days trend of the PMC_TRAN.

Then we manually updated those numbers by using ezcawrite in order not to reboot C1PSL.

So now it nicely indicates green in the medm screen.

 

  3027   Tue Jun 1 18:39:59 2010 NancyUpdate Lead spheres for the seismographs

 

the lead spheres that were placed below the granite slab have been flattened by hammering to have lesser degree of wobbling of the slab.

the height of each piece, and the flatness of their surfaces was checked by placing another slab over them and checking by the spirit level.

P6010170.JPG

P6010166.JPG

P6010164.JPG

 

  3026   Tue Jun 1 16:29:51 2010 AlbertoUpdateIOOMC transmitted beam aligned to the Faraday; next things to do

We moved the MC-trans pick-off mirror (= the beam splitter between the input of the Faraday and the steering mirror located right after MC3). Now the beam goes through the Farady without getting clipped.

This is the list of the things that have to be done next:

  1. take pictures of the beam spot just before and after the Faraday
  2. lock down to the table the MCTrans pickoff mirror with its screws
  3. measure the beam profile after the first MC telescope mirror (MMT1)
  4. remove Jenne's extra steering mirror from the MC table
  5. re-level the MC table with the bubble level
  6. align the MC-trans beam to its photodiode on the PSL table
  7. align the REFL beam to its photodiode on the AP table
  3025   Tue Jun 1 15:51:42 2010 steveUpdatePEMfoam box for good thermal stability

This box was made  to provide good thermal stability for seismometer calibration. There is an inner solid shell of 0.064" Al box that is covered by 2" insulation

inside and outside. The polystyrene foam is "CertiFoam 25 SE " from McMasterCarr #9255K3.

Attachment 1: albx2.PDF
albx2.PDF
  3024   Tue Jun 1 11:47:14 2010 steveUpdatePEMlead balls on concrete

Quote:

Valera and I put the 2 Guralps and the Ranger onto the big granite slab and then put the new big yellow foam box on top of it.

There is a problem with the setup. I believe that the lead balls under the slab are not sitting right. We need to cut out the tile so the thing sits directly on some steel inserts.

You can see from the dataviewer trend that the horizontal directions got a lot noisier as soon as we put the things on the slab.

 The tiles were cut out in 1.5" ID circle to insure that the 7/16" OD lead balls would not touch the tiles on Wednesday, May 26, 2010

Granite surface plate specifications: grade B, 18" x 24" x 3" , 139 lbs

These balls and granite plate were removed by  Rana in entry log #3018 at 5-31-2010

Attachment 1: P1060269.JPG
P1060269.JPG
  3023   Tue Jun 1 06:30:38 2010 KojiConfiguration40m Upgradingeffect of the arm length

I checked the effect of the arm length to the reflectance of the f2(=5*f1) sidebands.

Conclusion: If we choose L_arm = 38.4 [m], it looks sufficiently being away from the resonance
We may want to incorporate small change of the recycling cavity lengths so that we can compensate the phase deviation from -180deg.


f1 of 11.065399MHz is assumed. The carrier is assumed to be locked at the resonance.

Attachment 1: (Left) Amplitude reflectance of the arm cavity at f2 a a function of L_arm. (Right) Phase
Horizontal axis: Arm length in meter, Vertical Magnitude and Phase of the reflectance

At L=37.93 [m], f2 sidebands become resonant to the arm cavity. Otherwise, the beam will not be resonant.

Attachment 2: close-up at around 5 f1 frequency.
The phase deviation from the true anti resonance is ~0.7deg. This can be compensated by both PRC and SRC lengths.

Attachment 1: arm_cavity1.png
arm_cavity1.png
Attachment 2: arm_cavity2.png
arm_cavity2.png
Attachment 3: arm_cavity.nb.zip
  3022   Mon May 31 22:52:57 2010 ranaUpdatePEMGranite slab w/ lead balls is so far a flop

This plot shows the noise with the box on, but no granite. We're still pretty far off from the Guralp data sheet.

Untitled.png

I implemented software rotation in the huddle subtraction as Valera suggested and it works much better. The two plots below show the before and after. So far this is just 2 deg. of rotation around the z-axis. I'm assuming that aligning the seismometers vertically via bubble level is good enough for the z-axis, but I haven't calibrated the bubble yet.

huddlez.pnghuddlez.png

The residual slope is now suspiciously smooth. I somehow suspect that our readout electronics can still be responsible. We need to hook up a 9V battery to the input terminals to check it out. Its a little steeper than 1/f and I thought that we had exonerated the Guralp breakout box in the past, but now I'm not so sure. I'll let Jenne comment on that.

I also noticed that we have not yet divided by sqrt(2) to account for the fact that we are subtracting 2 seismometers. In principle, an unbiased estimate of the single seismometer noise will be lower by sqrt(2) than the green curve.

  3021   Mon May 31 17:47:34 2010 kiwamuUpdateIOOtoday's plan : MC alignment

[Alberto, Kiwamu]

0. have a coffee and then dress up the clean coat.

1. level the MC table

2. lock and align MC 

3. run A2L script to see how much off-centering of the spots

4. steer the periscope mirror <--- We are here

5. move the pick off mirror which is used for monitoring of MCT CCD

6. check the leveling and move some weights if it's necessary

7. shut down

  3020   Mon May 31 03:38:48 2010 KojiUpdateIOOMC alignment

Remember that you only can introduce the axis translations from the PSL table.
It is quite difficult to adjust the axis rotation.

The calibration factor from A2L results to the beam position is dx = (A2L_result - 1) *10.8mm

If I believer the result below, the spot positions on the mirrors are

MC1 Pitch      -1.1mm
MC1 Yaw        -0.20mm
MC3 Pitch      -1.5mm
MC3 Yaw        +0.35mm

This means that the beam is 1.3mm too high and 0.28mm too much in north

This corresponds to tilting SM2 by
0.33mrad in pitch (23deg in CW)
and
0.10mrad in yaw (7deg in CW).

Quote:

C1:SUS-MC1_ULPIT_GAIN =  0.900445
C1:SUS-MC1_ULYAW_GAIN =  0.981212
C1:SUS-MC3_ULPIT_GAIN =  0.86398

C1:SUS-MC3_ULYAW_GAIN =  1.03221  

 

Attachment 1: MC_spot_centering.png
MC_spot_centering.png
  3019   Mon May 31 00:10:18 2010 kiwamuUpdateIOOMC alignment

  [Alberto, Kiwamu]

The MC alignment is getting better by steering the axis of the incident beam into the MC.

We found the beam spot on MC1 and MC3 were quite off-centered in the beginning of today's work. It had the coil gain ratio of 0.6:1.4 after running the A2L script.

In order to let the beam hit the center of the MC1 and MC3, we steered the bottom mirror attached on the periscope on the PSL table to the yaw direction.

And then we got better numbers for the coil gain ratio (see the numbers listed at the bottom).

For the pitch direction, there still are some rooms to improve because we didn't do anything with the pitch. It is going to be improved tomorrow or later.

 

Here are the amounts of off-centering on MC1 and MC3 after steering the axis. 

 C1:SUS- MC1_ULPIT_GAIN =  0.900445

C1:SUS-MC1_ULYAW_GAIN =  0.981212

C1:SUS-MC3_ULPIT_GAIN =  0.86398

C1:SUS-MC3_ULYAW_GAIN =   1.03221

  3018   Sun May 30 22:18:49 2010 ranaUpdatePEMGranite slab w/ lead balls is so far a flop

The seismometers showed an increased noise in the Y-direction when put on top of the granite slab. By tapping the slab, you can tell that its really a mechanical resonance of the lead balls + granite system at ~15-20 Hz.

I tried new balls, flipping the slab upside down, and sitting on the slab for awhile. None of this changed the qualitative behavior, although each of the actions changed the resonance frequencies by several Hz.

I have removed the granite/balls and put the seismometers back on the linoleum floor. The excess noise is gone. I have put the new big box back on top of them and we'll see how the data looks overnight.

 

I expect that we should remove the linoleum in a wider area and put the seismometers directly on the floor.

  3017   Sun May 30 17:51:04 2010 kiwamuHowToPEMAllegra dataviewer

I found the dataviewer didn't work only on Allegra. This thing sometimes happened as described in the past entry.

I rebooted Allegra, then the problem was fixed.

 

  3016   Sun May 30 15:36:22 2010 AlbertoConfigurationPSLIMC periscope shutter

Two days ago I opened the PSL shutter by switching the switch on the shutter driver. That caused the shutter's switch on the medm screen to work in reversed mode: open meant closed and closed meant open.

I fixed that. Now the medm screen switch state is correct.

  3015   Sun May 30 15:33:21 2010 AlbertoConfigurationIOOmode cleaner and air conditioning

The mode cleaner is locked and the air conditioning is full on. So the the air conditioning doesn't seem to be so important for the lock to hold.

  3014   Sun May 30 13:26:07 2010 rana, kiwamuUpdatePSLnew HIGH-LOW value for PMC_TRANS

We changed the HIGH/LOW values of the PMC_TRANS.

The edited file was updated on the svn.


Since the PMC_TRANSPD was replaced behind the pzt mirror (see the entry), its nominal value were reduced to something like ~1V from the previous value of ~2V.

In the medm screen C1PSL_PMC.adl the PMC_TRAN always indicated red because the value were low compared with the previous one.

We went to /cvs/cds/caltech/target/c1psl, then edited psl.db

- Here are the new parameters we set up in the file.

grecord(ai,"C1:PSL-PMC_PMCTRANSPD")
{

  field(LOW,"0.98")
  field(LOLO,"0.93")
  field(HIGH,"1.15")
  field(HIHI,"1.3")

}

- - - -

These values are based on ~4days trend of the PMC_TRAN.

Then we manually updated those numbers by using ezcawrite in order not to reboot C1PSL.

So now it nicely indicates green in the medm screen.

  3013   Fri May 28 23:21:52 2010 KojiUpdateIOOMC alignment

Hm... You touched the optics between the MC and the Faraday... This will lead us to the painful work.

I am afraid that the beam is already walking off from the center of MC1/MC3 after the work on the PSL table.
This may result in the shift of the spot on those MC mirrors. So I recommend that:

- Lock the cavity
- Check the A2L for MC1/3
- Adjust it by the periscope
- If it is fine, adjust the optics after the MC (steering, Faraday, etc)

Off-centering of the MC2 spot is no problem. We can move it easily using Zach's scripts.
Tell me when the work is planed on Sunday as I might be able to join the work if it is in the evening.

Quote:

[Alberto, Kiwamu, Kevin, Rana]

Today we tried to measured the beam shape after the MC MMT1 that Jenne installed on the BS table.

The beam scan showed a clipped spot. We tracked it down to the Farady and the MCT pickoff mirror.

The beam was getting clipped at the exit of the Faraday. But it was also clipping the edge of the MCT pick-off mirror. I moved the mirror.

Also the beam looked off-center on MC2.

We're coming back on Sunday to keep working on this.

Now things are bad.

 

  3012   Fri May 28 21:32:32 2010 AlbertoUpdate40m UpgradingMC alignment

[Alberto, Kiwamu, Kevin, Rana]

Today we tried to measured the beam shape after the MC MMT1 that Jenne installed on the BS table.

The beam scan showed a clipped spot. We tracked it down to the Farady and the MCT pickoff mirror.

The beam was getting clipped at the exit of the Faraday. But it was also clipping the edge of the MCT pick-off mirror. I moved the mirror.

Also the beam looked off-center on MC2.

We're coming back on Sunday to keep working on this.

Now things are bad.

  3011   Fri May 28 14:24:30 2010 JenneUpdateVACvacuum work

I just got off the phone with Alberto and Kiwamu, and I'm going to try to recalculate things based on their measurements of the distances between MC3 and SM1.  It sounds like the CAD drawings we have aren't totally correct.    I know that when we opened doors just before Christmas we measured the distances between the BS table and the ITM tables, but I don't think we measured the distance between the IOO table and the BS table.  Hopefully we can fit everything in our chambers.....

  3010   Fri May 28 13:51:51 2010 JenneUpdateVACvacuum work

Quote:

We started a vacuum work in this morning. And still it's going on.

 

Although the last night the green team replaced a steering mirror by an 80% reflector on the PLS table, the beam axis to the MC looks fine.

The MC refl beam successfully goes into the MCrefl PD, and we can see the MC flashing as usual.

We started measuring the distance of the optics inside the vacuum chamber, found the distance from MC3 to MMT1(curved mirror) is ~13cm shorter than the design.

We moved the positions of the flat mirror after the Faraday and the MMT1, but could not track the beam very well because we did not completely lock the MC.

Now we are trying to get the lock of the MC by steering the MC mirrors.

 

 

 Hey guys,

I just finished redoing the calc based on the measurements that happened last week.  Using the average of the Vert and Horz measurements in Kevin's elog 2986, I find that we need to make the MMT telescope ~8cm longer.  So, can you please place the flat mirror after the Faraday in the same place as the drawing, but move the MMT1 79mm farther away from that flat mirror?  Looking at the table layouts that Koji has on the wiki, this should still (barely) fit.

New distances:

d2a = 884.0mm (no change) ------  MC3 to Flat after Faraday

d2b = 1123.2mm (move MMT1 farther toward center of BS table)  -------- Flat after Faraday (SM1) to MMT1

d3 = 1955.0mm (result of moving MMT1)  ---------  MMT1 to MMT2

d4a = 1007.9mm (no chnage)        ----------- MMT2 to SM2

d4b = 495.6mm (no change) ------------ SM2 to PRM

  3009   Fri May 28 13:32:01 2010 alberto, kiwamuUpdateVACvacuum work

We started a vacuum work in this morning. And still it's going on.

 

Although the last night the green team replaced a steering mirror by an 80% reflector on the PLS table, the beam axis to the MC looks fine.

The MC refl beam successfully goes into the MCrefl PD, and we can see the MC flashing as usual.

We started measuring the distance of the optics inside the vacuum chamber, found the distance from MC3 to MMT1(curved mirror) is ~13cm shorter than the design.

We moved the positions of the flat mirror after the Faraday and the MMT1, but could not track the beam very well because we did not completely lock the MC.

Now we are trying to get the lock of the MC by steering the MC mirrors.

 

P.S.

Kevin suceeded in locking it !!

  3008   Fri May 28 13:17:05 2010 josephbUpdateCDSFixed problem with channel access on c1iscex

Talked with Alex and tracked down why the codes were not working on the new c1iscex finally.  The .bashrc and .cshrc files in /home/controls/ on c1iscex has the following lines:

setenv EPICS_CA_ADDR_LIST 131.215.113.255
setenv EPICS_CA_AUTO_ADDR_LIST NO

This was interfering with channel access and preventing read and writes from working properly.  We simply commented them out. After logging out and back in, the things like ezcaread and write started working, and we were able to get the models passing data back and forth.

Next up, testing RFM communications between megatron on c1iscex.  To do this, I'd like to move Megatron down to 1Y3, and setup a firewall for it and c1iscex so I can test the frame builder and testpoints at the same time on both machines.

  3007   Fri May 28 11:35:33 2010 josephbUpdateCDSTaking a step backwards to get stuff running

I've modified the lsc.mdl and lsp.mdl files back to an older configuration, where we do not use an IO processor.  This seems to let things work for the time being on megatron while I try to figure out what the is wrong with the "correct" setup which includes the IO processor.

Basically I removed the adcSlave = 1 line in the cdsParameters block.

I've attached a screen shot of the desktop showing one filter bank in the LSP model passing its output correctly to a filter block in the LSC.  I also put in a quick test filter (an integrator) and you can see it got to 80 before I turned off the offset.

So far this is only running on megatron, not the new machine in the new Y end.

The models being use for this are located in /cvs/cds/caltech/cds/advLigoRTS/src/epics/simLink

Attachment 1: working_lsc_lsp.png
working_lsc_lsp.png
  3006   Fri May 28 11:26:35 2010 JenneUpdatePEMNew Foam Box installed

Quote:

Valera and I put the 2 Guralps and the Ranger onto the big granite slab and then put the new big yellow foam box on top of it.

There is a problem with the setup. I believe that the lead balls under the slab are not sitting right. We need to cut out the tile so the thing sits directly on some steel inserts.

You can see from the dataviewer trend that the horizontal directions got a lot noisier as soon as we put the things on the slab.

 You'll have to ask Steve how deep he cut, but the tile is cut around the lead balls, so they are not sitting on the linoleum.  They might just be sitting on the concrete slab, or whatever Steve found underneath the tile, instead of fancy steel inserts, but at least they're not on the tile.  I don't know why things got noisier though...

  3005   Fri May 28 10:44:47 2010 josephbUpdatePEMDAQ down

Quote:

 Although trends are available, I am unable to get any full data from in the past (using DTT or DV). I started the FB's daqd process a few times, but no luck. 

I blame Joe's SimPlant monkeying from earlier today for lack of a better candidate. I checked and the frames are actually on the FB disk, so its something else.

 I tried running dataviewer and dtt this morning.  Dataviewer seemed to be working.  I was able to get trends, full data on a 2k channel (seismic channels) and full data on a 16k channel (C1:PEM-AUDIO_MIC1)  This was tried for a period 24 hours a go for a 10 minute stretch.

I also tried dtt and was able to get 2k and 16k channel data, for example C1:PEM-AUDIO_MIC1.  Was this problem fixed by someone last night or did time somehow fix it?

  3004   Fri May 28 07:13:05 2010 AlbertoFrogsGreen LockingSR785 found abandoned next to the workbenches

A poor lonely SR785 was found this morning roaming around in the lab in evident violation of the fundamental rule which requires all the equipment on carts to be brought back inside the lab right after use.

The people and the professors related to the case should take immediate action to repair for their misdeed.

  3003   Fri May 28 00:40:53 2010 ranaUpdatePEMDAQ down

 Although trends are available, I am unable to get any full data from in the past (using DTT or DV). I started the FB's daqd process a few times, but no luck. 

I blame Joe's SimPlant monkeying from earlier today for lack of a better candidate. I checked and the frames are actually on the FB disk, so its something else.

  3002   Thu May 27 23:59:54 2010 ranaUpdatePEMNew Foam Box installed

Valera and I put the 2 Guralps and the Ranger onto the big granite slab and then put the new big yellow foam box on top of it.

There is a problem with the setup. I believe that the lead balls under the slab are not sitting right. We need to cut out the tile so the thing sits directly on some steel inserts.

You can see from the dataviewer trend that the horizontal directions got a lot noisier as soon as we put the things on the slab.

Attachment 1: Untitled.png
Untitled.png
  3001   Thu May 27 12:52:02 2010 AlbertoUpdate40m UpgradingArm lengths


For both sidebands to be antiresonant in the arms, the first modulation frequency has to be:

f1 = (n + 1/2) c / (2*L)

where L is the arm length and c the speed of light.  For L=38m, we pick to cases: n=3,  then f1a = 13.806231 MHz;  n=2, then f1b = 9.861594 MHz.

If we go for f1a, then the mode cleaner half length has to change to 10.857m.  If we go for f1b, the MC length goes to 15.200m. A 2 meter change from the current length either way.

And the mode cleaner would only be the first of a long list of things that would have to change. Then it would be the turn of the recycling cavities.

Kind of a big deal.

  3000   Thu May 27 10:30:32 2010 kiwamuHowToGreen LockingPSL setup for green locking

 I leave notes about a plan for the green locking especially on the PSL table.

 

 


 (1) open the door  of the MC13 tank to make the PSL beam go into the MC.  Lock it and then optimize the alignment of the MC mirror so that we can later align the incident beam from the PSL by using the MC as a reference.   

 (2) Remove a steering mirror located just after the PMC on the PSL table. Don't take its mount, just take only the optic in order not to change the alignment .

 (3) Put an 80% partial reflector on that mount to pick off ~200mW for the doubling . One can find the reflector on my desk.

 (4) Put some steering mirrors to guide the transmitted beam through the reflector to the doubling crystal. Any beam path is fine if it does not disturb any other setups. The position of the oven+crystal should not be changed so much, I mean the current position looks good.

 (5) Match the mode to the crystal by putting some lenses. The optimum conversion efficiency can be achieved with beam waist of w0~50um (as explained on #2735). 

 (6) Align the oven by using the kinematic mount. It takes a while. The position of the waist should be 6.7 mm away from the center of the crystal (as explained on #2850). The temperature controller for the oven can be found in one of the plastic box for the green stuff. After the alignment, a green beam will show up.

(8) Find the optimum temperature which gives the best conversion efficiency and measure the efficiency.

(7)  Align the axis of the PSL beam to the MC by steering the two mirrors attached on the periscope.

  2999   Thu May 27 09:43:50 2010 ranaUpdateGreen Lockingmore details

Quote:

 RMS which is integrated down to 1Hz  is 1.6MHz.

This number is almost what I expected assuming the cavity swings with displacement of x ~< 1um.

 Its OK, but the real number comes from measuring the time series of this in the daytime (not the spectrum). What we care about is the peak-peak value of the PZT feedback signal measured on a scope for ~30 seconds. You can save the scope trace as a PNG.

  2998   Thu May 27 08:22:57 2010 AidanUpdateComputersRestarted the elog this morning
  2997   Thu May 27 02:22:24 2010 kiwamuUpdateGreen Lockingmore details

 Here are some more plots and pictures about the end PDH locking with the green beam. 

-- DC reflection

 I expected that the fluctuation of the DC reflection had 1% from the resonant state to the anti-resonant state due to its very low finesse.

This values are calculated from the reflectivity of ETM measured by Mott before (see the wiki).

In my measurement I obtained  DC reflection of V_max=1.42 , V_min=1.30  at just after the PD.

These numbers correspond to 7.1% fluctuation. It's bigger than the expectation.

I am not sure about the reason, but it might happen by the angular motion of test masses (?)

 

--- time series

Here is a time series plot. It starts from openloop state (i.e. feedback disconnected).

At t=0 sec I connected a cable which goes to the laser pzt, so now the loop is closed.

You can see the DC reflection slightly decreased and stayed lower after the connection.

The bottom plot represents the feedback signal measured before a sum amp. which directly drives the pzt.

stimes.png

 

 

-- length fluctuation  

One of the important quantities in the green locking scheme is the length fluctuation of the cavity.

It gives us how much the frequency of the green beam can be stabilized by the cavity. And finally it will determine the difficulty of PLL with the PSL.

I measured a spectrum of the pzt driving voltage [V/Hz1/2] and then converted it to a frequency spectrum [Hz/Hz1/2].

I used the actuation efficiency of 1MHz/V for the calibration, this number is based on the past measurement.

spectrum.png

RMS which is integrated down to 1Hz  is 1.6MHz.

This number is almost what I expected assuming the cavity swings with displacement of x ~< 1um.

 

-- flashing

A picture below is a ETMx CCD monitor.

One of the spot red circled in the picture blinks when it's unlocked. And once we get the lock the spot stays bright.

ETMX_small.png

 

  2996   Wed May 26 22:22:03 2010 AlbertoConfiguration40m UpgradingArm cavity length

The second sideband is resonant in the arms for a cavity length of 37.9299m.

The nearest antiresonant arm lengths for f2 (55MHz) are 36.5753m and 39.2845m.

If we don't touch the ITMs, and we use the room we still have now on the end tables, we can get to 37.5m.

This is how the power spectrum at REFL would look like for perfect antiresonance:

reflRFpowerVsArmLength_3658m.png

And this is how it looks like for 37.5m:

reflRFpowerVsArmLength_3750m.png

Or, god forbid, we change the modulation frequencies...

  2995   Wed May 26 18:54:55 2010 AidanSummaryGreen LockingMounted Crystal 724 in the Doubling Oven

Andri and I mounted the Raicol Crystal #724 in one of the new Covesion Ovens. The procedure was the same as before - see elog entry here.

There was one issue - the glass plate that goes on top of the crystal is coated on one side with ITO (Indium-Tin Oxide) and it's not 100% certain that this was mounted in the correct orientation. It is virtually impossible to tell which side of the glass is coated.

The base plate of the oven was tapped for an M3 hole. We retapped it for an 8-32 and bolted it to a post and that one of the New Focus 4-axis translation stage. The assembly is currently bolted to the PSL table, awaiting use.

ELOG V3.1.3-